A man used his prototype invention, the 3D Maneuver Gear, to navigate his hometown. He maintained his velocity of 96 m/s as he swung on top of a 15-storey building, his anchors placed 250 m away, at the height of 325 m. He swung 52.43° to propel himself up, then 35° to the West to reach his endpoint. How fast is he in all 3D components?


1
Expert's answer
2022-03-10T18:01:26-0500

"v_x=\sqrt{2g(h-l)}\cos52.43°=23.4~\frac ms,"

"v_y= \n\\sqrt{2gl}\n\u200b\n \\sin35\u00b0=40.2 ~\n\\frac ms\n\u200b\n ,"

"v_z=v=96~\frac ms."